0 2 The product of Eigen values of the matrix P is P=4 -3 3 0 2 -1 (A) -6 (B) 2 (C) 6 (D) -2
Answer #1

P, .4.23 3 ام 2 : 6 == م -۴| تا 6 ? A ا .0.2- و 3 - 3 = 0 4 . مان ه: [ - 14+ 5- [، - (4) (مهران) + (1 - 2) 0 = 8 + (- 43 ار 1